Calculate the change in temperature of an 900mm aluminium bar, that has increased in length to 901.035mm.

 α = 25 x 10-6k-1

1
Expert's answer
2022-02-20T08:20:52-0500

Given:

"l_0=900\:\rm mm"

"\alpha=25*10^{-6}\:\rm K^{-1}"

"l=901.035\:\rm mm"


The thermal expansion is described by an equation


"\Delta l=l_0\alpha\Delta t""\Delta t=\frac{\Delta l}{\alpha l_0}\\
=\rm \frac{901.035-900}{25*10^{-6}*900}=46\:{^\circ C}"